40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 17 12:32:35 2021, Koji, Update, SUS, Remaining task for 2021 
    Reply  Fri Dec 17 19:19:42 2021, Koji, Update, SUS, Remaining task for 2021 PXL_20211218_030145356.MP.jpg
       Reply  Sun Dec 19 07:52:51 2021, Anchal, Update, SUS, Remaining task for 2021 
    Reply  Wed Dec 22 18:16:23 2021, Koji, Update, SUS, Remaining task for 2021 ITMX_2016_07_21.jpgScreen_Shot_2021-12-22_at_18.03.42.png
Message ID: 16519     Entry time: Fri Dec 17 12:32:35 2021     Reply to this: 16522   16534
Author: Koji 
Type: Update 
Category: SUS 
Subject: Remaining task for 2021 

Anything else? Feel free to edit this entry.

- SUS: AS1 hanging

- SUS: PR3/SR2/LO2 3/4" thick optic hanging

v Electronics chain check (From DAC to the end of the in-air cable / From the end of the in-air cable to the ADC)
[ELOG 16522]

- Long cable installation (4x 70ft)

- In-air cable connection to the flange

- In-vac wiring (connecting LO1 OSEMs)

- LO1 OSEM insertion/alignment

- LO1 Damping test

 

ELOG V3.1.3-